OpenStudy (anonymous):

It says to graph each linear equation how would i do this: x-2y=6

OpenStudy (anonymous):

Im sorry i still could'nt get it..I need major help

OpenStudy (anonymous):

solve for y

OpenStudy (anonymous):

I do but i end up getting it wrong, I have the answer in the back of the book but cant get it rite. Okay this is what im doin : x-2y=6, Then i subtrat from both side Which gives me -2y=6-x, Then i dive all of it by -2 which then gives me -3+1/2.. So what am i doing wrong..

OpenStudy (anonymous):

y=(1/2)x-3, 3= y intercept M or slope= (1/2)

OpenStudy (anonymous):

How would i do this one: x-y=-2

OpenStudy (anonymous):

-3=y intercept

OpenStudy (anonymous):

same thing solve for y, subtract x and divide by -1

OpenStudy (anonymous):

Where did you get the -3 from?

OpenStudy (anonymous):

sorry -3 was the first question

OpenStudy (anonymous):

Ya im also talking about the first question, How did you get -3?

OpenStudy (anonymous):

when you divide -2 from -2y=6-x you get y= -3- (-X/2) expand the negative to (x/2) making it a positive and rearange so positive is in front of the negative three

OpenStudy (anonymous):

Wait doesnt the -2 become a positive 2, because you need to get rid of the -2, rite.

OpenStudy (anonymous):

no you need to move the negative two from the y side to the x side